OpenStudy (anonymous):

Which of the following options is an equivalent function to f(x) = 2(5)^2x? choices: f(x) = 50^x f(x) = 100^x f(x) = 2(25)^x f(x) = 4(25)^x

OpenStudy (anonymous):

@green_1

OpenStudy (anonymous):

@Adi3

OpenStudy (green_1):

@Michele_Laino @Zale101

OpenStudy (adi3):

f(x) = 100^x

OpenStudy (anonymous):

can u help me with another? @Adi3

OpenStudy (adi3):

i just told you the answer

OpenStudy (anonymous):

Identify the domain of the exponential function shown in the following graph: choices: all real numbers x greater than or equal to 0 x greater than or equal to 0, 30000 is greater than or equal to x x greater than or equal to 7, 30000 is greater than or equal to x

OpenStudy (green_1):

i think it is A or B

OpenStudy (anonymous):

@bruno101

OpenStudy (anonymous):

@TylerD

OpenStudy (anonymous):

@green_1

OpenStudy (anonymous):

@baller4life

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@mathmate

OpenStudy (anonymous):

@TheSmartOne

TheSmartOne (thesmartone):

I see mathmate is typing a reply, so I'll leave it to him :)

OpenStudy (mathmate):

Hint for the original question: \(f(x)=a(b^{cx})=a[(bc)^x]\) by the law of exponents.

OpenStudy (mathmate):

Domain is the set of all valid input for the function, i.e. the valid values along the x-axis. For the question of domain, observe from the graph the valid values along the x-axis. Decide if x<0 is part of the domain, if x=0 is part of the domain, if x=10000 is part of the domain.

OpenStudy (anonymous):

so b

OpenStudy (anonymous):

@mathmate

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (mathmate):

I will be glad to comment on your answer if you would post the answer itself and not the assumed letter choice.

OpenStudy (anonymous):

x greater than or equal to 0 @mathmate

OpenStudy (mathmate):

Exactly. "x greater than or equal to 0" is the domain of the given function because the function starts at 0 towards the right, with no limitation where it stops. Also there are no vertical asymptotes or discontinuities.

OpenStudy (anonymous):

do you know the second one. @mathmate

OpenStudy (mathmate):

I thought this is the answer to the second question. Let me check.

OpenStudy (mathmate):

Do you mean the third one? Hint: average rate = \(\Large \frac{f(x_2)-f(x_1)}{x_2-x_1}\) where in your case x1=1, x2=4.

OpenStudy (anonymous):

yeah i know the formula and i got -55/3 but the reasoning behind that answer make no sense @mathmate

OpenStudy (mathmate):

Can you show me how you found -55/3? We'll try to fix it if it does not work! :)

OpenStudy (anonymous):

-14/3 is what i meant @mathmate

OpenStudy (anonymous):

@TylerD

OpenStudy (mathmate):

Yes, -14/3 is correct, obtained as (2-16)/3=-14/3. Well done!